Airlines usually over-book the seats on an aircraft by a certain margin because they know from experience that some people change or do not show for their scheduled flight. Data collected for a particular Melbourne–Darwin flight showed that, on average, 183 people (with a standard deviation of 31) did arrive for their scheduled flight. The data followed a normal distribution. The aircraft has seats for 344 passengers.
a) A particular flight goes every day. During one year of operation, how many times would you expect there to be more passengers than available seats?
b) The cost of each flight to the airline is $71000. The cost of a ticket is $460. On what proportion of flights does the airline lose money.
c) In a period of 12 days, what is the probability the airline loses money?
I have tried that the first question ,but my answer is 1, I am so confusing about that and also the following questions .THX a lot.

Answer by ewatrrr(24785)


Hi
a)with z = (344-183)/31 = 5.1935....
how many times would you expect there to be more passengers than available seats? never
b) $71000/$460 = 154
z = (154-183)/31 = - .9243
P(z<-.9245) = 17.77% of the time
c) .1777*12 = 2.1324 Approximately 2 days
my thoughts...
Wish You the Best in your Studies.